# A Survey on Personalized Alignment—The Missing Piece for Large Language Models in Real-World Applications

### v3

@@ 1 Introduction
Figure 2: Overview of personalized alignment framework.
2 Preliminaries
@@ 2.3 From Social to Personalized Preference
Agent Workflow Customization Xie et al. (2024); Wang et al. (2024d, f); Guan et al. (2024) , yellowbox ] ] ] [Alignment through Feedback (§4.3),bluebox [Training-time Alignment,bluebox [Standard Alignment Jang et al. (2023); Park et al. (2024), Parameter-Efficient Alignment QI et al. (2024); Tan et al. (2024b); Hebert et al. (2024), Agent Alignment Guan et al. (2024), Federated Alignment Huang et al. (2024b, c); Kuang et al. (2024); Fan et al. (2024),yellowbox ] ] [Inference-time Alignment,bluebox [ Reward-guided Decoding Chen et al. (2025), Prediction Ensemble Shi et al. (2024), Output Refinement Yang et al. (2024a), yellowbox ] ] [Multi-objective Optimization,bluebox [ Direct Objective Space Optimization Li et al. (2024c), Preference-guided Optimization Guo et al. (2024); Wang et al. (2024c), Preference Learning through Interaction Shao et al. (2024); Wang et al. (2024b), yellowbox ] ] ] ]
@@
$\pi_{\rm s}(y|c,h)={\mathbb{E}_{u\in\mathcal{U}}}[\pi_{\rm p}(y|u,c,h)],$
where $\pi_{\rm s}$ represents the social preference distribution adopted as the learning objective in current alignment practice, $\pi_{\rm p}$ denotes the personalized preference distribution for individual $u\in\mathcal{U}$, and $h$ represents the universal human values.
@@ 3 Formalizing Personalized Alignment
Figure 2 illustrates our framework for personalized alignment (§3.1), guided by hierarchical alignment objectives (§3.2). Figure 3 presents an example.
@@ 4.1 Preference Memory Management
Personalization requires a dedicated memory mechanism to utilize user-specific knowledge across interactions Zhang (2023). This section examines preference types and methods for inferring preference from various sources.
Preference Types.
User preferences comprise explicit and implicit preferences Jawaheer et al. (2014); Zhang et al. (2024d); Li et al. (2025). Explicit preferences represent directly stated user attitudes and requirements, typically expressed as clear positive, negative, or neutral stances toward specific attributes or behaviors. Implicit preferences, in contrast, are derived from indirect signals that reveal user characteristics, which can be gathered from: (1) User-Generated Content: Textual data such as social media posts, reviews, and articles that reveal expertise levels, interests, and writing styles. (2) Demographic Attributes: Self-reported structured information, e.g., age, gender, and educational background. (3) Behavioral Signals: Interaction patterns captured through comparative judgments, ratings, click-through behaviors, etc. While these implicit signals don’t directly state preferences, they offer rich contextual information for inferring user characteristics and tendencies.
@@
Explicit preferences offer precise personalization signals but remain sparse in practice due to users’ reluctance, inability to articulate, or unawareness of their preferences Lee et al. (2024b). Implicit preferences, while information-rich, present noise, redundancy, and inconsistency challenges Preetha and Shankar (2014). These challenges motivate the need for preference inference - a process that distills explicit preferences by projecting implicit preferences onto predefined structured preference spaces Lee et al. (2024b). However, significant challenges remain in defining meaningful preference dimensions that capture user intent and handling temporal dynamics of preferences Liu (2015).
@@ 4.2 Personalized Generation and Rewarding
These methods represent preferences in latent spaces through: (1) contextual embeddings compressing user data into fixed representations Ning et al. (2024); Li et al. (2024e); Shenfeld et al. (2025), and (2) latent variable frameworks learning preference distributions Siththaranjan et al. (2024); Poddar et al. (2024); Gong et al. (2024), capturing uncertainty and multimodal patterns in user preferences. Both approaches balance preference preservation with computational efficiency.
@@ 4.3 Alignment through Feedback
Training-time alignment strategies vary across personalization approaches and model accessibility. When base models are accessible, standard SFT and RL optimization can be directly applied Park et al. (2024); Jang et al. (2023) for prompting-based, encoding-based, and full-parameter personalization. For adapter-based and agent-based approaches, training focuses only on user-specific components: adapter parameters QI et al. (2024); Tan et al. (2024b); Hebert et al. (2024) and agent modules Guan et al. (2024) respectively. When base models are inaccessible, personalized federated learning Huang et al. (2024c); Kuang et al. (2024); Fan et al. (2024) enables privacy-preserving distributed training through architectural innovations Yi et al. (2024); Saadati et al. (2025) and dynamic adaptation Bao et al. (2023b); Lee et al. (2024a).
@@
Recent advances achieve personalized alignment through inference-phase decoding modifications, avoiding costly training, including: (1) reward-guided decoding, where PAD Chen et al. (2025) generates token-level personalized rewards to steer model predictions; (2) prediction ensemble, where MOD Shi et al. (2024) and Drift Kim et al. (2025) combine token-level predictions from multiple objective-specific models; and (3) output refinement, where MetaAligner Yang et al. (2024a) introduces an external correction model to refine outputs towards desired objectives. While these approaches offer immediate adaptability and training-free operation, they face inherent trade-offs between real-time adaptability and long-term preference learning due to the lack of parameter updates.
@@ 4.4 Resources
Benchmark Size Feedback (Annotator) Preference Dimensions Evaluation Metrics Preference Type Task Universal Value Alignment TruthfulQA Lin et al. (2021) 817 Ranking (Human) Truthfulness Rating/Win/ACC/BLEU/ROUGE Explicit Preference Instruction-Following HHH-Alignment Askell et al. (2021) 221 Ranking (Human) Helpfulness, Honesty, Harmlessness ACC Explicit Preference Instruction-Following Self-Instruct Wang et al. (2022) 11,800 Ranking (Human&AI) Helpfulness Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following MT Bench Zheng et al. (2023) 3,360 Ranking (Human&AI) Helpfulness Rating/Win/ACC/BLEU/ROUGE Explicit Preference Instruction-Following Advbench Zou et al. (2023) 150 Ranking (AI) Harmlessness Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following AlpacaEval 2.0 Dubois et al. (2024) 805 No Feedback Helpfulness Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following Arena Hard Li et al. (2024d) 500 No Feedback Helpfulness Rating/Win/ACC/BLEU/ROUGE Explicit Preference Instruction-Following, Code Generation RewardBench Lambert et al. (2024) 2,985 Ranking (Human) Helpfulness ACC Explicit Preference Chat, Chat-Hard, Safety, Reasoning Personalized Alignment P-Soups Jang et al. (2023) 50 No Feedback Expertise, Informativeness, Style Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following LaMP Salemi et al. (2023) 25,095 No Feedback Not Explicitly Defined Rating/Win/BLEU/ROUGE User-Generated Content Text Classification, Text Generation OpinionQA Santurkar et al. (2023) 1,176 Ranking (Human) Not Explicitly Defined ACC Behavioral Signals Question-Answering GlobalOpinionQA Durmus et al. (2023) 2,556 Ranking (Human) Not Explicitly Defined ACC Behavioral Signals Question-Answering FLASK Ye et al. (2023) 1,740 No Feedback Not Explicitly Defined Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following REGEN Sayana et al. (2024) 403,000 Rating (Human) Not Explicitly Defined BLEU/ROUGE/Similarity Scores User-Generated Content, Behavioral Signals Conversational Recommendations LongLaMP Kumar et al. (2024a) 9,658 No Feedback Not Explicitly Defined ROUGE User-Generated Content, Email Completion, Abstract Generation, Demographic Attributes Review Writing, Topic Writing PGraphRAG Au et al. (2025) 10,000 Text, Rating (Human) Not Explicitly Defined Rating/Win/BLEU/ROUGE User-Generated Content, Behavioral Signals Long and Short Text Generation, Classification PersonalLLM Zollo et al. (2025) 10,000 No Feedback Not Explicitly Defined Rating/Win Behavioral Signals Instruction-Following ALOE Wu et al. (2025) 100 No Feedback Not Explicitly Defined Win Demographic Attributes Instruction-Following PERSONA Castricato et al. (2025) 3,868 Ranking (AI) Not Explicitly Defined Win/ACC Demographic Attributes Instruction-Following
@@ 8 Key Challenges and Future Directions
Foundational challenges focus on understanding and modeling user preferences: (1) capturing complex personalized preferences that evolve temporally Chandrashekaran et al. (1996), are influenced by social context Izuma and Adolphs (2013), vary across scenarios (e.g., work vs. life situations) Smailagic and Siewiorek (2002), and exhibit dynamic trade-offs between different dimensions (e.g., quality vs. speed) Li et al. (2024c); (2) addressing data-related challenges, including both the scarcity of high-quality personalization data (in terms of scale, diversity, and temporal dynamics) and the cold-start problem where systems lack initial user data for meaningful personalization Schein et al. (2002); (3) developing evaluation frameworks that can reliably assess personalization quality beyond surface-level metrics; (4) ensuring models understand user needs Zhang et al. (2023).
Technical challenges concern the implementation of personalized systems: (1) integrating multimodal signals (e.g., text, vision, audio) to better understand and generate personalized content Pi et al. (2024); (2) incorporating user preferences into long-chain reasoning while preserving interpretability and contextual consistency River Dong et al. (2025); (3) balancing performance and computational efficiency for real-time personalization; (4) ensuring personalized models can effectively generalize from simple to complex tasks with weak supervisions Burns et al. (2023); Kenton et al. (2024).

### v4

@@ 2.3 From Social to Personalized Preference
$\pi_{\rm s}(y|c)={\mathbb{E}_{u\in\mathcal{U}}}[\pi_{\rm p}(y|u,c)],$
where $\pi_{\rm s}$ represents the social preference distribution adopted as the learning objective in current alignment practice, $\pi_{\rm p}$ denotes the personalized preference distribution for individual $u\in\mathcal{U}$.
@@
Agent Workflow Customization Xie et al. (2024); Wang et al. (2024d, f); Guan et al. (2024) , yellowbox ] ] ] [Alignment through Feedback (§4.3),bluebox [Training-time Alignment,bluebox [Standard Alignment Jang et al. (2023); Park et al. (2024), Parameter-Efficient Alignment QI et al. (2024); Tan et al. (2024b); Hebert et al. (2024), Agent Alignment Guan et al. (2024), Federated Alignment Huang et al. (2024b, c); Kuang et al. (2024); Fan et al. (2024),yellowbox ] ] [Inference-time Alignment,bluebox [ Reward-guided Decoding Chen et al. (2025b), Prediction Ensemble Shi et al. (2024), Output Refinement Yang et al. (2024a), yellowbox ] ] [Multi-objective Optimization,bluebox [ Direct Objective Space Optimization Li et al. (2024c), Preference-guided Optimization Guo et al. (2024); Wang et al. (2024c), Preference Learning through Interaction Shao et al. (2024); Wang et al. (2024b), yellowbox ] ] ] ]
@@ 3 Formalizing Personalized Alignment
Figure 2 illustrates our framework for personalized alignment (§3.1), guided by hierarchical alignment objectives (§3.2).
@@ 4.1 Preference Memory Management
Personalization requires a dedicated memory mechanism to utilize user-specific knowledge across interactions Zhang (2023). In particular, the sparse distribution of user preferences in interactions poses significant challenges for retrieving relevant preferences from long contexts Zhao et al. (2025); Pan et al. (2025). Since the capacity of LLMs in processing long-form text is beyond our scope, we assume the memory contains only preference-relevant information sources. This section examines both explicit and implicit preferences Jawaheer et al. (2014); Zhang et al. (2024d); Li et al. (2025), and methods for inferring explicit preferences from various implicit sources.
Explicit Preferences.
Explicit preferences represent directly stated user stances toward specific attributes or behaviors (i.e., preference dimensions). Formally, consider a preference space with $D$ dimensions $\{d_{1},d_{2},\cdots,d_{D}\}$. The explicit preference of a user $u\in\mathcal{U}$ can be represented by a normalized stance vector $\boldsymbol{p}^{u}=[p^{u}_{1},p^{u}_{2},\cdots,p^{u}_{D}]$. Each $p^{u}_{i}\in[-1,1]$ refers to the stance value for the dimension $d_{i}$, where positive values indicate favorable attitudes, negative values indicate unfavorable attitudes, and zero indicates neutrality. The magnitude $|p^{u}_{i}|$ represents the strength of the stance. With these explicit preference representations, one can easily formalize how dimensional evaluations contribute to users’ overall judgment of model responses, enabling straightforward personalized generation and reward computation. Specifically, given a model response $y\in\mathcal{Y}$ in context $c\in\mathcal{C}$, the response is evaluated with a user-agnostic rating vector $\boldsymbol{r}(y|c)=[r_{1}(y|c),r_{2}(y|c),\cdots,r_{D}(y|c)]$ in the preference space, where each dimension is rated independently of user preferences. The overall judgment of $u$ for $y$ is assumed as the dot product of the explicit preference vector and rating vector [^1] The linear combination modeling is enough to capture complex preference patterns through appropriate dimension transformations (e.g., non-linear mappings). To improve interpretability and optimization efficiency, we suggest constructing orthogonal dimensions through variable substitution and linear transformation (e.g., Principal Component Analysis) when needed.:
$R^{u}(y|c)=(\boldsymbol{p}^{u})^{\mathrm{T}}\boldsymbol{r}(y|c)$
This formulation enables direct optimization of personalized generation models by maximizing the expected overall judgment. Notably, the preference vector $\boldsymbol{p}^{u}$ may depend on both the context $c$ and dimensional ratings $\boldsymbol{r}$. For example, a user might prefer professional responses in technical discussions but favor empathetic ones in casual conversations. Furthermore, even in the same context, when a response receives high ratings in professionalism, the user’s preference for response length might shift from negative (preferring concise) to positive (favoring detailed), demonstrating how ratings in one dimension can influence preferences in another. However, due to implementation complexity, most existing studies assume explicit preferences remain constant regardless of context and model responses Zhao et al. (2025); Li et al. (2025).
Implicit preferences.
Implicit preferences refer to indirect signals that reveal user characteristics, vaguely reflecting users’ judgments on specific model responses. As exemplified in Figure 3, these signals can be observed through: (1) User-Generated Content: Textual data such as social media posts, chat records, reviews, and articles that exhibit expertise levels, interests, and writing styles. (2) Demographic Attributes: Self-reported structured information, e.g., age, gender, educational background, etc. (3) Behavioral Signals: Interaction patterns captured through comparative judgments, ratings, click-through behaviors, etc. While these implicit signals do not directly state preferences, they offer rich contextual information for inferring user characteristics and tendencies without explicitly defining preference dimensions. However, unlike explicit preferences that enable direct computation of judgments for unseen responses through preference-rating aggregation, implicit preferences require data-driven learning to map indirect signals to personalization objectives, such as response generation and reward computation, necessitating careful design of preference modeling algorithms.
@@
Explicit preferences offer precise personalization signals but remain sparse in practice due to users’ reluctance, inability to articulate, or unawareness of their preferences Lee et al. (2024b). Implicit preferences, while information-rich, present challenges in noise, redundancy, and inconsistency Preetha and Shankar (2014). Although end-to-end modeling from implicit preferences to personalized generation and rewarding is possible, such approaches often lack interpretability and control over the learned preference structure. These challenges motivate preference inference - a process that distills explicit preferences by projecting implicit preferences onto predefined structured preference spaces Lee et al. (2024b). Given several examples reflecting implicit preferences of a user, the inference process typically involves three steps: (1) constructing a structured preference space by identifying key dimensions that capture user preferences; (2) estimating the user’s explicit preferences on the preference space from individual examples, which are often partial, ambiguous, and noisy; and (3) aggregating potentially conflicting dimensional preferences across examples to obtain a robust unified explicit preference vector. Without considering contextual and rating effects, simple averaging of stance values suffices Li et al. (2025); otherwise, data-driven learning is needed to model how stance values vary with specific context and response rating vectors. Nevertheless, significant challenges remain in defining preference dimensions, as well as capturing the temporal evolution of user preferences Liu (2015).
@@ 4.2 Personalized Generation and Rewarding
These methods represent preferences in latent spaces through: (1) contextual embeddings compressing user data into fixed representations Ning et al. (2024); Li et al. (2024e); Shenfeld et al. (2025), and (2) latent variable frameworks learning preference distributions Siththaranjan et al. (2024); Poddar et al. (2024); Gong et al. (2024); Chen et al. (2025a), capturing uncertainty and multimodal patterns in user preferences. Both approaches balance preference preservation with computational efficiency.
@@ 4.3 Alignment through Feedback
Training-time alignment strategies vary across personalization approaches and model accessibility. When base models are accessible, standard SFT and RL optimization can be directly applied Park et al. (2024); Jang et al. (2023) for prompting-based, encoding-based, and full-parameter personalization. For adapter-based and agent-based approaches, training focuses only on user-specific components: adapter parameters QI et al. (2024); Tan et al. (2024b); Hebert et al. (2024) and agent modules Guan et al. (2024) respectively. When base models are inaccessible, personalized federated learning Huang et al. (2024c); Kuang et al. (2024); Fan et al. (2024) enables privacy-preserving distributed training through architectural innovations Yi et al. (2024); Saadati et al. (2025); Tran et al. (2025) and dynamic adaptation Bao et al. (2023b); Lee et al. (2024a).
@@
Recent advances achieve personalized alignment through inference-phase decoding modifications, avoiding costly training, including: (1) reward-guided decoding, where PAD Chen et al. (2025b), Amulet Zhang et al. (2025) and CoS He et al. (2025) generates token-level personalized adjustments to steer model predictions; (2) prediction ensemble, where MOD Shi et al. (2024) and Drift Kim et al. (2025) combine token-level predictions from multiple objective-specific models; and (3) output refinement, where MetaAligner Yang et al. (2024a) introduces an external correction model to refine outputs towards desired objectives. While these approaches offer immediate adaptability and training-free operation, they face inherent trade-offs between real-time adaptability and long-term preference learning due to the lack of parameter updates.
@@ 4.4 Resources
Benchmark Size Feedback (Annotator) Preference Dimensions Evaluation Metrics Preference Type Task Universal Value Alignment TruthfulQA Lin et al. (2021) 817 Ranking (Human) Truthfulness Rating/Win/ACC/BLEU/ROUGE Explicit Preference Instruction-Following HHH-Alignment Askell et al. (2021) 221 Ranking (Human) Helpfulness, Honesty, Harmlessness ACC Explicit Preference Instruction-Following Self-Instruct Wang et al. (2022) 11,800 Ranking (Human&AI) Helpfulness Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following MT Bench Zheng et al. (2023) 3,360 Ranking (Human&AI) Helpfulness Rating/Win/ACC/BLEU/ROUGE Explicit Preference Instruction-Following Advbench Zou et al. (2023) 150 Ranking (AI) Harmlessness Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following AlpacaEval 2.0 Dubois et al. (2024) 805 No Feedback Helpfulness Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following Arena Hard Li et al. (2024d) 500 No Feedback Helpfulness Rating/Win/ACC/BLEU/ROUGE Explicit Preference Instruction-Following, Code Generation RewardBench Lambert et al. (2024) 2,985 Ranking (Human) Helpfulness ACC Explicit Preference Chat, Chat-Hard, Safety, Reasoning Personalized Alignment P-Soups Jang et al. (2023) 50 No Feedback Expertise, Informativeness, Style Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following LaMP Salemi et al. (2023) 25,095 No Feedback Not Explicitly Defined Rating/Win/BLEU/ROUGE User-Generated Content Text Classification, Text Generation OpinionQA Santurkar et al. (2023) 1,176 Ranking (Human) Not Explicitly Defined ACC Behavioral Signals Question-Answering GlobalOpinionQA Durmus et al. (2023) 2,556 Ranking (Human) Not Explicitly Defined ACC Behavioral Signals Question-Answering FLASK Ye et al. (2023) 1,740 No Feedback Not Explicitly Defined Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following REGEN Sayana et al. (2024) 403,000 Rating (Human) Not Explicitly Defined BLEU/ROUGE/Similarity Scores User-Generated Content, Behavioral Signals Conversational Recommendations LongLaMP Kumar et al. (2024a) 9,658 No Feedback Not Explicitly Defined ROUGE User-Generated Content, Email Completion, Abstract Generation, Demographic Attributes Review Writing, Topic Writing PGraphRAG Au et al. (2025) 10,000 Text, Rating (Human) Not Explicitly Defined Rating/Win/BLEU/ROUGE User-Generated Content, Behavioral Signals Long and Short Text Generation, Classification PersonalLLM Zollo et al. (2025) 10,000 No Feedback Not Explicitly Defined Rating/Win Behavioral Signals Instruction-Following ALOE Wu et al. (2025) 100 No Feedback Not Explicitly Defined Win Demographic Attributes Instruction-Following PERSONA Castricato et al. (2025) 3,868 Ranking (AI) Not Explicitly Defined Win/ACC Demographic Attributes Instruction-Following PREFEVAL Zhao et al. (2025) 3,000 Ranking (Human) Not Explicitly Defined Rating/ACC Explicit Preference, User-Generated Instruction-Following Content, Behavioral Signals
@@ 8 Key Challenges and Future Directions
Foundational challenges focus on understanding and modeling user preferences: (1) capturing complex personalized preferences that evolve temporally Chandrashekaran et al. (1996), are influenced by social context Izuma and Adolphs (2013), vary across scenarios (e.g., work vs. life situations) Smailagic and Siewiorek (2002), and exhibit response-dependent trade-offs between different dimensions (e.g., quality vs. speed) Li et al. (2024c); (2) addressing data-related challenges, including both the scarcity of high-quality personalization data (in terms of scale, diversity, and temporal dynamics) and the cold-start problem where systems lack initial user data for meaningful personalization Schein et al. (2002); (3) developing evaluation frameworks that can reliably assess personalization quality beyond surface-level metrics; (4) ensuring models understand user needs and follow user preference Zhang et al. (2023); Zhao et al. (2025); (5) enabling proactive preference exploration and learning, where systems actively engage with users to discover and refine preference inference, and balance exploration of new preferences with exploitation of known preferences.
Technical challenges concern the implementation of personalized systems: (1) integrating multimodal signals (e.g., text, vision, audio) to better understand and generate personalized content Pi et al. (2024, 2025); (2) incorporating user preferences into long-chain reasoning while preserving interpretability and contextual consistency River Dong et al. (2025); (3) balancing performance and computational efficiency for real-time personalization; (4) ensuring personalized models can effectively generalize from simple to complex tasks with weak supervisions Burns et al. (2023); Kenton et al. (2024).